check whether the limit of the function f (x,y)=3x^3y/x^6+2y ^2 exists as (x,y) to(0,0)
1
Expert's answer
2020-03-16T12:11:51-0400

Let "f(x,y)=\dfrac{3x^3y}{x^6+2y^2}." Let's approach "(0,0)" along the "x-"axis.

Then "y=0" gives "f(x,0)=\dfrac{3x^3(0)}{x^6+2(0)^2}=0" for all "x\not=0."

"f(x,y)\to0" as "(x,y)\to(0,0)" along the "x-"axis.


Let's approach "(0,0)" along the line "y=x^3."

Then "y=x^3" gives

"f(x,x^3)=\dfrac{3x^3(x^3)}{x^6+2(x^3)^2}=1"

"f(x,y)\to1" as "(x,y)\to(0,0)" along the line "y=x^3."

Since "f" has two different limits along two different lines, the given limit


"\lim\limits_{(x,y)\to(0,0)}={3x^3y \over x^6+2y^2}"

does not exist.